OpenStudy (e.cociuba):

Solve by completing the square: 7b^2 - 14b - 56= 0

OpenStudy (mertsj):

Factor out 7 first.

OpenStudy (e.cociuba):

?

OpenStudy (mertsj):

??

OpenStudy (anonymous):

Dude, if your gonna play games, please don't bother(:

jimthompson5910 (jim_thompson5910):

\[\large 7b^2 - 14b - 56 = 0 \] \[\large 7(b^2 - 2b - 8) = 0 \] \[\large b^2 - 2b - 8 = \frac{0}{7} \] \[\large b^2 - 2b - 8 = 0 \] I'll let you continue

OpenStudy (e.cociuba):

Ok so this is where I got:  Ok so I need help of figuring out wat I have to do w/ the b^2...

jimthompson5910 (jim_thompson5910):

your next step is to take half of the b coefficient -2 -2/2 = -1 then you square this to get (-1)^2 = 1

jimthompson5910 (jim_thompson5910):

so you add 1 to both sides